STERLING TIRE C Fur the Year Ended December 31, 20X1 1,372,000 936,000 bl csts (26.000 ties al 335) Ficed cuss Eaminga batore Intanat ad axa nbarest apanse Eamings before texes EBT Income tax expense :30%; 423,000 $ 296,100 .Comaute the degree e opraning evrage (Round your answer to 2 decimal ploces.) b. Cupue lhe degtee ol Iancs levege. (Round your answer to 2 decimal places af fnancial Ieverage c. Compute the degrce of combinod leveagc Round your answer to 2 decimal places. d. Compure the break even point in units. Round your answer to the nearest whole number.) Hints References eBcok & Resources Check mu 103 PM E E O lype here to search
Answer #1

Operating Leverage = (Sales - Variable cost) / EBIT = (1872000 - 936000) / 476000 = 1.97

Financial Leverage = EBIT / EBT = 476000 / 423000 = 1.13

Combined Leverage = Operating Leverage x Financial Leverage = 1.97 x 1.13 = 2.23

Breakeven point in units = Fixed cost / contribution per unit = 460000 / (72 - 36 ) = 12778 units